| ID | 
Date | 
Author | 
Category | 
Subject | 
Text | 
  | 
| 
  79  
 | 
Wed Jun 26 11:59:12 2024 | 
Ruijiu, Shahab, Carsten | Analysis | Prepare plots for Thomas's presentation | data:
 
(env_OnlineDataAnalysisSystem) litv-exp@lxbk0497:/lustre/astrum/experiment_data/2024-05_E018/OnlineDataAnalysisSystem/data/iq/IQ_2024-06-
 
12_21-47-54$
 | 
 6x   | 
| 
  78  
 | 
Sat Jun 15 03:46:34 2024 | 
Night shift | Accelerator | GSI down | All quadrupole and dipole magnets at SIS and 
ESR are out of order. Repairs can be expected 
earliest tomorrow evening. No shifts will | 
   | 
| 
  77  
 | 
Fri Jun 14 14:44:51 2024 | 
Afternoon shift | Runs | Afternoon shift | Pizza shift
 
 
[B]the scan direction is from higher wavelength | 
   | 
| 
  76  
 | 
Fri Jun 14 13:14:29 2024 | 
Konstantin, Bernhard, Peter, Lukas | General | DAQ crashes at :10 | 10 Minutes past a full hour (XX:10) the windows 
PCs (sometimes) loose connection.
 
This causes the labview server & the laser | 
   | 
| 
  74  
 | 
Fri Jun 14 11:44:23 2024 | 
Carsten Brandau | Accelerator | Cooling Time in SC7 increased from 9s to 12s | Cooling Time in SC7 increased from 9s to 12s | 
   | 
| 
  73  
 | 
Fri Jun 14 06:38:41 2024 | 
Konstantin Mohr, Bernhard Maass, Peter Micke, Lukas Kau, Dounia Boudefla | Runs | Morning runs | Very successful shift, no crashes until 11:10
 
Previous night shift had problems with the 
pulse laser starting around 3am.
 | 
   | 
| 
  72  
 | 
Thu Jun 13 22:08:07 2024 | 
Dayshift | Accelerator | Problems with Accelerator? | It seems to be that there are some problems 
with the injections in the ESR.
 
The last 2 accumulation cycles showed that | 
      | 
| 
  71  
 | 
Thu Jun 13 18:48:26 2024 | 
Ruijiu Chen, David, Shahab, Yuri | Analysis | The ion identification of 229Th is correct. | Harmonic: 124.0 , Bp: 7.796046 [Tm]
 
ion            fre[Hz]                   
    yield [pps]    
 | 
     | 
| 
  70  
 | 
Thu Jun 13 18:38:24 2024 | 
Rodolfo Sanchez, Julian Palmes, Sebastian Klammes | Laser | Laser Dye change - new MRC settings - scraper checks of laser beam | We changed the dye at ~3pm.
 
 
After some fine adjustements of the laser | 
    | 
| 
  69  
 | 
Thu Jun 13 18:23:46 2024 | 
Julian | Laser | Restarting the laser controller | If the laser controller is locked up again, 
the little display will show an error message 
after an attempted medusa scan.
 | 
   | 
| 
  68  
 | 
Thu Jun 13 18:03:36 2024 | 
Afternoon shift | Runs | Continuation: First scans | [TABLE border="1"]
 
Time | Part. Det. Int | ESR-intensity| cooler-HV 
| LMD-DAQ1 Files | LMD-DAQ2 Files | scan | 
     | 
| 
  67  
 | 
Thu Jun 13 14:56:09 2024 | 
Stefan Schippers | Accelerator | Estimated beam lifetime: About  150 s | Estimate of storage lifetime using my code 
'hydrocal':
 
 | 
   | 
| 
  66  
 | 
Thu Jun 13 14:41:22 2024 | 
Ruijiu Chen, David | Analysis | Lifetime of beam during June 13 9:00~11:00 |   | 
    | 
| 
  65  
 | 
Thu Jun 13 14:35:27 2024 | 
Ruijiu Chen, David | Analysis | Lifetime of beam during June 13 7:00~9:00 |   | 
    | 
| 
  64  
 | 
Thu Jun 13 14:20:36 2024 | 
Ruijiu Chen, David | Analysis | Lifetime of beam during June 13 4:30:7:00 |   | 
    | 
| 
  63  
 | 
Thu Jun 13 13:29:47 2024 | 
Ruijiu Chen, David | Analysis | Lifetime of beam during June 12 22:00- June 13 0:00 |   | 
    | 
| 
  62  
 | 
Thu Jun 13 12:37:14 2024 | 
Ruijiu Chen, David | Analysis | Lifetime of beam in center of mass is 193 s. |   | 
    | 
| 
  61  
 | 
Thu Jun 13 12:05:31 2024 | 
Konstantin, Peter, Jonas | Accelerator | RF-amplitude | To realize the shortest bunches possible, 
we check the bunch length as a function of 
two different bunching amplitudes. With the | 
     | 
| 
  60  
 | 
Thu Jun 13 10:42:02 2024 | 
Danyal | General | Question | What is the ion beam lifetime, according to 
the DCCT and Schottky? 
 
Can somebody measure it, please? 
 | 
   | 
| 
  59  
 | 
Thu Jun 13 10:32:29 2024 | 
Konstantin | Accelerator | (Proper) Cooler settings | We change the acceleration voltage of the 
electron cooler and observe the platform 
high voltage and the count rate at the particle | 
   | 
| 
  58  
 | 
Thu Jun 13 10:06:35 2024 | 
Constantine, Peter, Lukas, Jonas, Dounia, Carsten | Accelerator | Suspected Beam Troubles (is it really 229-Thorium?) | From experience, the lifetime of the beam 
in the storage ring should be on the order 
of magnitude of hours, however, we observer | 
   | 
| 
  57  
 | 
Thu Jun 13 02:05:04 2024 | 
Carsten | Accelerator | Short Lifetime at 100 mA | Lifetime still is short | 
    | 
| 
  55  
 | 
Wed Jun 12 22:48:55 2024 | 
Carsten |   | Beam Halo | The seamingly short live time reported in 
[URL=https://elog.gsi.de/esr/E0052/53]entry 
53[/URL] is due to a halo from a different | 
   | 
| 
  54  
 | 
Wed Jun 12 21:21:15 2024 | 
Yuri | Detectors | NTCap started | Fcenter - 244.6 MHz
 
BW - 10 MHz | 
   | 
| 
  53  
 | 
Wed Jun 12 18:48:43 2024 | 
Afternoon shift | Runs | First laser scans | [TABLE border="1"]
 
Time | Part. Det. Int | ESR-intensity| cooler-HV 
| LMD-DAQ1 Files | LMD-DAQ2 Files | scan | 
    | 
| 
  51  
 | 
Wed Jun 12 17:12:13 2024 | 
Julian, Danyal, Rodolfo | Laser | Timing für 229Th89+ | We did the timing Ion-Beam Laser-beam time 
overlap. 
 
 | 
    | 
| 
  49  
 | 
Wed Jun 12 12:12:44 2024 | 
Konstantin Mohr, Julien Spahn, Jonas Koedel, Bernhard Maass | Accelerator | First Scraper Scan | Thorium is in the ESR, we have time to check 
the beam position with the scrapers.
 
Beam Intensity in the ring 6E-5
 | 
    | 
| 
  48  
 | 
Wed Jun 12 09:43:57 2024 | 
Carsten | Accelerator | Stopped NTCAP at 9:43 | Setup at ESR is carried on - stopped the long-time 
NTCAP run from overnight, | 
   | 
| 
  47  
 | 
Wed Jun 12 09:17:43 2024 | 
Carsten |   | Tests of background with ionization pumps on (from monday) | On monday the ionization pumps were switch 
on to test the background, and if it is possible 
to run with the pumps switched on.
 | 
       | 
| 
  46  
 | 
Wed Jun 12 01:51:51 2024 | 
RuiJiu, David | Analysis | ID | The preliminary identification has been thoroughly 
checked using the RionID software. We have 
confirmed that the identification is accurate.
 | 
       | 
| 
  45  
 | 
Tue Jun 11 23:46:06 2024 | 
Carsten )))) | Accelerator | ID | ----
 
The frequency of the primary beam 238U82+ 
at the injection:
 | 
   | 
| 
  44  
 | 
Tue Jun 11 22:59:06 2024 | 
Carsten | Accelerator | Identification / NTCAP started | ESR is set-up in most parts.We are trying 
to identify our 229Th89+. It is more cumbersome 
then exspected because of the thick target | 
   | 
| 
  43  
 | 
Tue Jun 11 21:38:35 2024 | 
Julian, Imke, Rodolfo | Laser | Wavemeter HeNe long time stability | The HeNe was running over night and the measured 
laser frequency of the wavemeter was logged, 
see picture.
 | 
    | 
| 
  42  
 | 
Tue Jun 11 21:23:11 2024 | 
Julian, Imke | General | How to read out voltage | 1.) If not already opened, open the Medusa 
Laser Control software, see picture:
 
2.) In the "view" submenu, choose "proteus | 
      | 
| 
  41  
 | 
Tue Jun 11 09:17:46 2024 | 
Julien, Konstantin, Lukas | DAQ | Important commands (DAQ1) | To connect to the DAQ1 pc, open a terminal 
and type:
 
 ssh atplaser@r4l-41
 | 
        | 
| 
  40  
 | 
Mon Jun 10 11:59:12 2024 | 
Ruijiu Chen | Analysis | The ion identification of spectrum recorded on 2022 | The simulated frequency is given in simulation_result.out. 
The yield is calculated by LISE file. The 
LISE file is attached here also. | 
      | 
| 
  39  
 | 
Mon Jun 10 09:17:38 2024 | 
Carsten | DAQ | Updated Channel List |   | 
     | 
| 
  38  
 | 
Mon Jun 10 08:57:35 2024 | 
Carsten/Konstantin | DAQ | Further Tests from Sunday 2024-06-09 - new voltages and tresholds for PMTs - testdata in files | Voltages and thresholds increased for PMT 
Middle and North (in order to account for 
the ADC measurement in DAQ2
 | 
   | 
| 
  37  
 | 
Sun Jun  9 10:20:38 2024 | 
Carsten/Danyal | Detectors | Recap from yesterdays work  with test beam (Detectors / beam /DAQ) | In the ESR we had thw whole day (saturday) 
a 400MeV/u bunched 238U92+ test beam (about 
2 MHz). The RF signal is cabled as input | 
   | 
| 
  36  
 | 
Sun Jun  9 10:08:24 2024 | 
Carsten/Danyal | Detectors | Particle detectors - Recap of yesterdays work | [B]South (behind cooler) - "SA_SZ" - DD2...[/B]
 
We have put the particle detecotr North to 
the poistion behind the cooler (north) to | 
   | 
| 
  35  
 | 
Sat Jun  8 13:11:52 2024 | 
Carsten | General | Put the go4login in atplaser .bashrc | Put the go4login in atplaser .bashrc in order 
to avoid that one always has to go to the 
go4 directory to set the environment variable.
 | 
   | 
| 
  34  
 | 
Sat Jun  8 12:45:05 2024 | 
Carsten | DAQ | Temporary change of RF signal | The setting was 1.540974 MHz and division 
factor 68 (for a 200 MeV/u) beam, and the 
signal was taken from a signal generator.
 | 
   | 
| 
  33  
 | 
Sat Jun  8 12:02:20 2024 | 
Carsten | Detectors | Test Particle Detectors - Det South not working | With the test beam from ESR I started to have 
a look at the PartDets.:
 
North outside: Works properly with a voltage | 
   | 
| 
  32  
 | 
Fri Jun  7 17:10:45 2024 | 
Sebastian | General | XUV LED | Maximum XUV LED Voltage: 6.5V | 
   | 
| 
  31  
 | 
Fri Jun  7 15:29:50 2024 | 
Rodolfo | Detectors | Stored beam @ ESR | One of the first PMT signals from the stored-ion 
beam @ ESR | 
    | 
| 
  30  
 | 
Fri Jun  7 11:14:04 2024 | 
Rodolfo | Laser | Trigger-signal for MRC | We are using a 2-Channel Function Generator 
(from Textronik) to generate the trigger 
signal for the MRC Beam Stabilization. 
 | 
       | 
| 
  29  
 | 
Fri Jun  7 09:15:31 2024 | 
Carsten | DAQ | Updated Channel List |   | 
     | 
| 
  28  
 | 
Thu Jun  6 19:51:11 2024 | 
Carsten | Accelerator | ESR crew left after persisting problems with ESR magnets | No beam today.--. | 
   | 
| 
  27  
 | 
Thu Jun  6 19:47:01 2024 | 
Carsten | DAQ | Exchange of Mesytec CFD/shaper - noise on signal lines from ESR magnets | Jan exchanged the Mesytec CFD/shper since 
it could not be "remote controlled" anymore. 
The new one has different shaping times and | 
   | 
| 
  26  
 | 
Thu Jun  6 19:42:20 2024 | 
Carsten | Detectors | Exchange of amplifer channels / cables channels / HV voltage cables for PMT south | In order to check if the problem of regular 
burst of pulses on the PMT south is related 
to amplfier channels and or the HV cable, | 
   | 
| 
  25  
 | 
Thu Jun  6 09:39:23 2024 | 
Carsten | DAQ | Preliminary Channel Table (DAQ1 and DAQ2) |   | 
     | 
| 
  24  
 | 
Thu Jun  6 09:23:53 2024 | 
Konstantin Mohr | DAQ | Timing | The best way to adjust the bunch-laser timing 
is to start go4 (mbs must be running) from 
atppc023. Therefore use x2goclient and use | 
      | 
| 
  22  
 | 
Thu Jun  6 08:16:36 2024 | 
Danyal | Accelerator | scraper info by Robert Boywitt | attched is more info about the T-scrapers 
(E-cooler) by Robert Boywitt (Tel: 2067 or 
3357) | 
       | 
| 
  Draft  
 | 
Thu Jun  6 07:54:15 2024 | 
Danyal | General | pressure gauges and ion pumps  | yesterday (05.06.2024), firstly, the two ion 
pumps (underneath the fluorescence chamber) 
were switched off by the vacuum group (Björn | 
   | 
| 
  20  
 | 
Thu Jun  6 07:12:49 2024 | 
Konstantin Mohr | Laser | T-Scraper | The T-Scrapers can be driven by DeviceControl, 
which must be started from a console. To 
do so, first open the app-launcher and then | 
 6x   | 
| 
  19  
 | 
Wed Jun  5 19:22:43 2024 | 
Carsten, Konstantin, Rodolfo | Detectors | PMT tests and preliminary thresholds | For all channels of the CAEN CFD the width 
of the logic signal is set to 255ns (maximum) 
to mitigate potential double pulses. For | 
     | 
| 
  18  
 | 
Wed Jun  5 14:17:36 2024 | 
Carsten | Accelerator | Distance Cooler Detector | Calculation from Helmut Weick (copied from 
E142)
 
Distance e-cooler to particle detectors chamber:
 | 
   | 
| 
  17  
 | 
Tue Jun  4 22:54:02 2024 | 
Rodolfo | Detectors | PMT - Manufacture - Test Result Data |   | 
      | 
| 
  16  
 | 
Mon Jun  3 19:42:06 2024 | 
Rodolfo | Laser | Laser-beam through E-Cooler | UV-laser beam is now going through the Electron-Cooler. 
 
 | 
    | 
| 
  15  
 | 
Mon Jun  3 18:24:44 2024 | 
Rodolfo | Detectors | PMT arrangement | The picture shows the actual configuration 
of the PMTs.
 
 | 
     | 
| 
  14  
 | 
Sun Jun  2 00:01:46 2024 | 
Carsten | DAQ | 25ps TDC CAEN V1290N DAQ2 lead to crash in DAQ2 due to a hardware issue | On the CAEN V1290N TDC the piggy board for 
the lower 8 channels is not fixed well to 
the motherboard.
 | 
   | 
| 
  13  
 | 
Sat Jun  1 19:19:07 2024 | 
Carsten | Detectors | Reminder: Check signals in scope and in DAQ of detectors with voltages switched off (vs. CFD threshold) | When the ion pumps and vacuum gauges are switched 
off the scaler data of 
 
PMT_south, PMT_middle, PMT_north, PD_NO, | 
   | 
| 
  12  
 | 
Sat Jun  1 19:00:57 2024 | 
Carsten | DAQ | Change of (offline) frequency for common stop (CS) of  VUPROMs (and 2nd TDC) and scaling factor | Temporarily, for tests of the second DAQ I 
changed the frequency from the frequency 
generator and scaling factor  that is used | 
   | 
| 
  11  
 | 
Sat Jun  1 18:55:18 2024 | 
Carsten | Analysis | Go4 Analysis Version for DAQ1 (atplaser) | Small enhancements and minor corrections / 
bug fixes are made to the original version 
of the online GO$ for DAQ1.
 | 
   | 
| 
  10  
 | 
Sat Jun  1 18:14:01 2024 | 
Carsten | Analysis | DAQ and analysis directories | [U][B]DAQ1 (original):[/B][/U]
 
-user atplaser
 
-runs on lxg1297
 | 
   | 
| 
  9  
 | 
Sat Jun  1 18:00:20 2024 | 
Carsten | DAQ | ReadOut Trigger im Timer Scaler changed because of faulty signal from frequency generator | Old readout trigger (~200 Hz) in 1MHz Scaler 
showed two nearby peaks indicating a jitter 
of the readout frequency or same faulty output | 
     | 
| 
  8  
 | 
Fri May 31 16:49:52 2024 | 
Carsten | DAQ | Wiring DAQ 1 | Started to Check the wiring.
 
 
Status as of  2024-05-31
 | 
   | 
| 
  7  
 | 
Fri May 31 15:53:06 2024 | 
Ruijiu Chen | DAQ | run NTCAP system | NTCAP is working again. 
 
Connect the RF cable and the scaler cables 
to NTCAP system.
 | 
   | 
| 
  6  
 | 
Wed May 29 10:54:21 2024 | 
Shahab, Jan | General | Gas Target Events with new timing system | Using the new timing system, it is possible 
to grab any events as pulses, but also as 
long inter event gates under GENESYS. Many | 
       | 
| 
  5  
 | 
Wed May 29 09:49:30 2024 | 
Rodolfo | Laser | Laser-beam profile | Using a telescope (Plan-concave lens -500 
mm and Plan-convex lens +1000 mm) after the 
pulsed-dye laser system we got the following | 
     | 
| 
  4  
 | 
Mon May 27 13:37:31 2024 | 
Jan | DAQ | MesyTec Shaper Control | For the 2nd DAQ the MesyTec Shaper ([URL=https://elog.gsi.de/esr/NucAR+lab/230418_110112/MesyTec_MSCF16-F-V.pdf]MSCF 
manual[/URL]) is used as a spectroscopy amplifier 
and CFD.
 | 
     | 
| 
  3  
 | 
Thu Apr 25 16:34:05 2024 | 
Rodolfo | Laser | Efficiency of Fluorescein27 | The figure below shows the efficiency of Fluorescein27 
vs. Wavelength. Each point was averaged over 
500 laser shots.
 | 
    | 
| 
  2  
 | 
Tue Apr 16 16:28:21 2024 | 
Danyal | General | Great | Thanks Jan! | 
   | 
| 
  1  
 | 
Tue Apr 16 14:42:37 2024 | 
Jan | General | G-PAC Proposal |   | 
    |